# OBsmith: LLM-powered JavaScript Obfuscator Testing

### v1

@@ 3.2. LLM-powered Sketch Generation
OBsmith generates sketches—i.e., program templates with typed placeholders that capture domain knowledge about JavaScript—by prompting an LLM with a constrained template language and explicit formatting requirements (Figure 3). The goal is to obtain (i) diverse yet valid templates that exercise representative control-flow and data-flow patterns, and (ii) sketches that can be deterministically instantiated by our program generator in §3.5.
@@
The prompt (Figure 5 in appendix) specifies: (1) the sketching DSL (literal, reference and expression placeholders); (2) usage constraints (operands may be literals, references, or nested expressions); (3) an I/O format that yields exactly ten sketches in a fenced code block; and (4) a worked example that pairs a template with a filled instance to ground the LLM’s understanding of intended semantics. This mix of schema, constraints, and exemplars reduces drifting and encourages syntactically uniform outputs that downstream tooling can parse.
@@
To broaden coverage, the prompt explicitly requests common JavaScript idioms and control-flow constructs (loops, branches, function calls, array operations), while leaving randomness via placeholders. OBsmith further promotes heterogeneity through: (1) multi-sketch requests (ten distinct templates per invocation); (2) operator sets in expression placeholders (e.g., arithmetic(…, +, -, *, /)), from which the filler later samples; and (3) nesting of expression generators to produce deeper ASTs. Together, these encourage varied CFG/DFG shapes from a single, concise specification.
@@
Figure 4. Program generation workflow (left) and differential testing with ground truth workflow (right)
@@ 3.3. Feedback Loop for Sketch Generation
While LLMs can produce diverse and syntactically valid sketches from carefully designed prompts (§3.2), their output may drift toward common or uninformative patterns. At the same time, differential testing (§3.4) often produces valuable signals in the form of failures, inconsistencies, or execution traces that highlight semantic corner cases. The feedback loop is designed to exploit these signals: instead of discarding failures after bug discovery, OBsmith repurposes them to guide future sketch generation.
@@
The feedback loop works synergistically with differential testing (§3.6) and metamorphic testing (§3.7). While testing reveals reusable patterns from failing programs, the feedback loop ensures that LLMs actively explore related semantic neighborhoods. This combined mechanism balances data-driven generalization (via LLM) and deterministic reuse (via testing), enabling both novelty and stability in the evolving sketch corpus.
@@ 3.4. Sketch Extraction from Real-World Program
Differential testing and metamorphic testing (§3.6, §3.7) frequently reveal program fragments that expose semantic inconsistencies in obfuscators. These fragments often encode subtle interactions between JavaScript features—such as scoping rules, exception propagation, or dynamic evaluation—that are difficult to anticipate in manual prompt design. Automatically converting such fragments into parameterized sketches increases test diversity while focusing future iterations on bug-revealing structures. Moreover, sketch extraction can automatically use existing JavaScript programs, which usually contain the developer’s domain knowledge and have been proven to be effective for compiler testing(Zang et al., 2024b). For the above two reasons, OBsmith also designed an automated framework to extract sketches from existing programs.
@@
OBsmith implements the extraction in AST level, it parses the concreate JavaScript program to a Babel AST and proceeds the sketch extraction in two steps:
@@ 2. Step 2:
    Replacement. The program is parsed into an Abstract Syntax Tree (AST), and literals, variable references, and selected expressions are abstracted back into OBsmith placeholders (numberLiteral, booleanReference, arithmetic(), etc.). The generalization step ensures that the resulting sketch is type-correct and can be re-instantiated with diverse values.expressions
@@
Together with LLM-powered sketch generation, program filling, and differential testing, sketch extraction completes a feedback-driven workflow that continually enriches OBsmith’s test suite. This synergy ensures that OBsmith is not only capable of finding bugs once, but also of systematically amplifying its effectiveness over time by reusing and refining bug-inducing patterns.
@@ 3.5. Sketch-based Program Generation
OBsmith generate sketches using LLMs and auto-extraction, these sketches are input to the program generator. Program generation is a two-step process that transforms partial sketches into complete, executable JavaScript programs which are ready for differential and metamorphic testing.
@@ 3.5.1. Sketch filling
The sketch solving is a specialized code transformation phase to transform an abstract sketch into multiple, unique, and concrete JavaScript programs. It systematically fills the literal and reference placeholders with random value or variables and uses an ”expression generator” to fill expression placeholders within a sketch, effectively translating the high-level program template into executable JavaScript code. This process is the first and critical step in creating the diverse corpus of programs required for effective differential testing.
@@
(numberLiteral and booleanLiteral). These are the simplest transformations, typically involving the replacement of an Identifier node in the AST. (numberLiteral, booleanLiteral): When an identifier matching one of these placeholders is found, it is replaced with a new Literal node in the AST. The value of this literal is a randomly generated primitive of the corresponding type. For booleanLiteral, OBsmith simply replace it with a True or a False with equal probability. Since JavaScript only provides the Number class and doesn’t distinguish between Integer and Float numbers, for numberLiteral, OBsmith generates an random integer or a random float with equal probability.
@@
It’s worth noting that in addition to using explicitly defined variables (let a = 1; let b = True), OBsmith also supports filling sketches with elements from arrays. Unlike other programming languages that contains strict type check mechanism, JavaScript allows elements of different types to exist in the same array. For example, let arr = [True, 3.1415926, "Hello World!", 1]; is a valid statement in JavaScript. If arr is accessible in current scope, OBsmith can use arr[0] to fill booleanReference and use arr[1] or arr[3] to fill numberReference.
@@
(e.g. arithmatic(operand1, operand2, operator1, operator2 …)) This is the most sophisticated part of the sketch filling algorithm. The expression placeholders are identified as CallExpression nodes in Babel AST. When a CallExpression node whose callee is one of the generators (arithmetic, relation, logic) is found, the script performs the following steps:
@@ 3.5.2. Program Enhancing
Program enhancement is a source-to-source transformation that takes a filled JavaScript program (in §3.5.1)as input and systematically injects logging code to produce an enhanced program. The goal is to make the program’s execution process observable and deterministic but not change it’s behavior and functionality. OBsmith implements the program enhancement in Babel. The input (filled) program is first parsed into an AST, which OBsmith traverses top-down using a visitor pattern. During traversal, new ExpressionStatement nodes are injected at locations such as BlockStatement or FunctionDeclaration, with each ExpressionStatement node containing a CallExpression (e.g., console.log) for logging. After traversal, the enhanced AST is converted back to JavaScript code and written to an output file. The enhancement involves several different types of logging code injection, each designed to provide different aspects of runtime observability. Specifically, we inject the following logging instruments.
@@
3.6. Differential Testing
Differential testing is a widely adopted technique in compiler validation (Yang et al., 2011; Zang et al., 2023), where the absence of a reliable ground truth necessitates comparing the outputs of multiple compiler variants on the same input program. In contrast, OBsmith operates under the assumption that the original JavaScript program, executed directly in a trusted engine (Node.js), serves as the definitive ground truth. This allows OBsmith to adapt differential testing specifically for the obfuscation domain.
@@ 3.6. Differential Testing
By leveraging the original program as an oracle, OBsmith strengthens differential testing with a reliable ground truth. This adaptation enables precise detection of semantic inconsistencies introduced by obfuscation, distinguishing it from traditional approaches that rely on consensus among potentially flawed variants.
@@ 3.7. Metamorphic Testing
Differential testing (§3.6) is our primary mechanism for checking semantic equivalence between an original program and its obfuscated counterpart. In addition, we explore whether metamorphic testing can stress obfuscators by applying equivalence-preserving rewrites to the inputs before obfuscation and then verifying that obfuscation does not change program behavior. The goal here is not limited to the “single-obfuscator available” setting; rather, it is to probe whether obfuscators are robust to semantics-preserving variation in their inputs.
@@
OBsmith defines a suite of metamorphic relations (MRs) over JavaScript programs, our implementation use three most frequently used MRs in compiler testing:
@@
By encoding program equivalences as metamorphic relations, OBsmith provides a lightweight yet powerful mechanism to validate obfuscators in the absence of oracles. This technique, in combination with differential testing, yields comprehensive correctness checking across both multi-obfuscator and single-obfuscator scenarios.
@@ 4.1.1. Large language models
Regarding the choice of LLM, we chose the most powerful version of the LLM model currently on the market. For each llm, we generate 100 different sketches. We also test the coding agent and see its effectiveness in generating sketchs compared to directly use base model. For fair comparison since coding agent like Gemini-CLI or Claude Code can see the whole project, we generate sketches first and then put other LLMs results into our project directory. Specifically, we use the following models: Gemini 2.5 Pro, GPT-5, Claude 4 Opus, Qwen3-235B-A22B-2507, Grok 4. Besides, we also evaluate LLM agent’s ability in sketch generation, we use Gemini CLI with Gemini 2.5 Pro to generate 100 sketches as well. In order to fair comparison, we begin with Gemini CLI without other model’s result, in other word, we do this first.
@@ 4.1.2. Obfuscators under test
Based on Google’s investigation(Jiang et al., 2025), Obfuscator.IO(obi, 2024) and JS-Confuser are two most wide used JavaScript obfuscator in Google’s production. So our main targets are the two most popular JavaScript obfuscators: JS-Confuser and Obfuscator.IO. JavaScript obfuscation takes candidate program as input alongside a JSON configuration.
Recognizing that the behavior of an obfuscator can vary significantly based on its settings, we systematically applies different configurations to each candidate program. This results in the generation of multiple, distinctively obfuscated versions of each original program, thereby maximizing the coverage of the obfuscator’s feature set and potential transformation space. JS-Confuser have 3 preset configurations - low, medium and high. We didn’t make any changes to its configurations. Obfuscator.IO have 4 preset configurations - Default, Low, Medium, and High. We use these preset.
However, our OBsmith implementation use console log to check program equivalence, but in the low, medium, and high settings, the console was disabled so we change the option to false. For high level, Obfuscator.IO enable debug protection which will make the debug tools useless but make the script hangs forever, so we also set that option to false. We didn’t make any other changes on configurations. Compact (put all things in one line) is also disabled in setting since V8 will reject code with too long line trigger a javascript invalid size error.

### v2

@@ 3.2. LLM-powered Sketch Generation
OBsmith generates sketches—i.e., program templates with typed placeholders that capture domain knowledge about JavaScript—by prompting an LLM with a constrained template language and explicit formatting requirements (Fig. 4). The goal is to obtain (i) diverse yet valid templates that exercise representative control-flow and data-flow patterns, and (ii) sketches that can be deterministically instantiated by our program generator in §3.5.
@@
The prompt (Fig. 5 in appendix) specifies: (1) the sketching DSL (literal, reference and expression placeholders); (2) usage constraints (operands may be literals, references, or nested expressions); (3) an I/O format that yields exactly ten sketches in a fenced code block; and (4) a worked example that pairs a template with a filled instance to ground the LLM’s understanding of intended semantics. This mix of schema, constraints, and exemplars reduces drifting and encourages syntactically uniform outputs that downstream tooling can parse.
@@
We ensure diversity at two stages: sketch synthesis and sketch instantiation. (1) Sketch synthesis (LLM stage). Firstly, OBsmith’s prompt explicitly asks for representative JavaScript idioms and control-flow constructs (e.g., loops, branches, function calls, and array operations) and requests distinct sketches per invocation, encouraging structural variety rather than repeated patterns. Furthermore, we generate sketches in a single LLM session, keeping prior outputs in context and explicitly prompting the model to avoid producing sketches similar to earlier ones.
(2) Sketch instantiation (filling stage). OBsmith then expands each sketch into many concrete programs by sampling at controlled variability points: literals are randomized; references are selected via scope analysis from in-scope variables of the appropriate type; and expression factories sample an operator from an explicit operator set (and may nest recursively to increase AST depth). Finally, we repeat this filling procedure multiple times per sketch, producing a set of distinct programs from the same template.
Together, these mechanisms yield diverse CFG/DFG shapes and expression structures from a compact sketch specification.
@@
Figure 4. Program generation workflow (left) and reference-oriented equivalence testing workflow (right)
@@ 3.3. Feedback Loop for Sketch Generation
While LLMs can produce diverse and syntactically valid sketches from carefully designed prompts (§3.2), their output may drift toward common or uninformative patterns. At the same time, reference-oriented equivalence testing (§3.7) often produces valuable signals in the form of failures, inconsistencies, or execution traces that highlight semantic corner cases. The feedback loop is designed to exploit these signals: instead of discarding failures after bug discovery, OBsmith repurposes them to guide future sketch generation.
@@
The feedback loop works synergistically with reference-oriented equivalence testing (§3.7) and metamorphic testing (§3.8). While testing reveals reusable patterns from failing programs, the feedback loop ensures that LLMs actively explore related semantic neighborhoods. This combined mechanism balances data-driven generalization (via LLM) and deterministic reuse (via testing), enabling both novelty and stability in the evolving sketch corpus.
@@ 3.4. Sketch Extraction from Real-World Program
Reference-oriented equivalence testing and metamorphic testing (§3.7, §3.8) frequently reveal program fragments that expose semantic inconsistencies in obfuscators. These fragments often encode subtle interactions between JavaScript features—such as scoping rules, exception propagation, or dynamic evaluation—that are difficult to anticipate in manual prompt design. Automatically converting such fragments into parameterized sketches increases test diversity while focusing future iterations on bug-revealing structures. Moreover, sketch extraction can automatically use existing JavaScript programs, which usually contain the developer’s domain knowledge and have been proven to be effective for compiler testing (Zang et al., 2024b). For the above two reasons, OBsmith also designed an automated framework to extract sketches from existing programs.
@@
OBsmith implements the extraction at the AST level. It parses the concrete JavaScript program to a Babel AST and proceeds the sketch extraction in two steps:
@@ 2. Step 2:
    Replacement. The program is parsed into an Abstract Syntax Tree (AST), and literals, variable references, and selected expressions are abstracted back into OBsmith placeholders (numberLiteral, booleanReference, arithmetic(), etc.). The generalization step ensures that the resulting sketch is type-correct and can be re-instantiated with diverse values and expressions.
@@
Together with LLM-powered sketch generation, program filling, and reference-oriented equivalence testing, sketch extraction completes a feedback-driven workflow that continually enriches OBsmith’s test suite. This synergy ensures that OBsmith is not only capable of finding bugs once, but also of systematically amplifying its effectiveness over time by reusing and refining bug-inducing patterns.
@@ 3.5. Sketch-based Program Generation
OBsmith generates sketches using LLMs and auto-extraction, and these sketches are input to the program generator. Program generation is a two-step process that transforms partial sketches into complete, executable JavaScript programs which are ready for equivalence checking.
@@ 3.5.1. Sketch filling
The sketch filling is a specialized code transformation phase to transform an abstract sketch into multiple, unique, and concrete JavaScript programs. It systematically fills the literal and reference placeholders with random values or variables and uses an ”expression generator” to fill expression placeholders within a sketch, effectively translating the high-level program template into executable JavaScript code. This process is the first and critical step in creating the diverse corpus of programs required for effective testing.
@@
(numberLiteral and booleanLiteral). These are the simplest transformations, typically involving the replacement of an Identifier node in the AST. (numberLiteral, booleanLiteral): When an identifier matching one of these placeholders is found, it is replaced with a new Literal node in the AST. The value of this literal is a randomly generated primitive of the corresponding type. For booleanLiteral, OBsmith simply replaces it with a True or a False with equal probability. Since JavaScript only provides the Number class and doesn’t distinguish between Integer and Float numbers, for numberLiteral, OBsmith generates a random integer or a random float with equal probability.
@@
It’s worth noting that in addition to using explicitly defined variables (let a = 1; let b = True), OBsmith also supports filling sketches with elements from arrays. Unlike other programming languages that contain strict type check mechanism, JavaScript allows elements of different types to exist in the same array. For example, let arr = [True, 3.1415926, "Hello World!", 1]; is a valid statement in JavaScript. If arr is accessible in current scope, OBsmith can use arr[0] to fill booleanReference and use arr[1] or arr[3] to fill numberReference.
@@
(e.g. arithmetic(operand1, operand2, operator1, operator2 …)) This is the most sophisticated part of the sketch filling algorithm. The expression placeholders are identified as CallExpression nodes in Babel AST. When a CallExpression node whose callee is one of the generators (arithmetic, relation, logic) is found, the script performs the following steps:
@@ 3.5.2. Program Enhancing
Program enhancement is a source-to-source transformation that takes a filled JavaScript program (in §3.5.1)as input and systematically injects logging code to produce an enhanced program. The goal is to make the program’s execution process observable and deterministic but not change its behavior and functionality. OBsmith implements the program enhancement in Babel. The input (filled) program is first parsed into an AST, which OBsmith traverses top-down using a visitor pattern. During traversal, new ExpressionStatement nodes are injected at locations such as BlockStatement or FunctionDeclaration, with each ExpressionStatement node containing a CallExpression (e.g., console.log) for logging. After traversal, the enhanced AST is converted back to JavaScript code and written to an output file. The enhancement involves several different types of logging code injection, each designed to provide different aspects of runtime observability. Specifically, we inject the following logging instruments.
@@
3.6. Correctness Criteria and Observational Equivalence
Goal and scope.
An obfuscator is intended to be semantics-preserving: the obfuscated program should exhibit the same behavior as the original one when executed in the same runtime environment (e.g., a fixed Node.js version). Because fully proving semantic equivalence for JavaScript is infeasible in general, OBsmith adopts an observational equivalence criterion based on a normalized observable trace that is (i) substantially stronger than comparing only final outputs, and (ii) directly checkable at scale.
Observable trace.
Given a program $P$ and an input $x$ (OBsmith programs may take no explicit input; $x$ may be empty), we execute $P$ under a fixed environment $E$ with OBsmith instrumentation enabled and obtain a sequence of runtime events:
$\mathsf{Trace}_{E}(P,x)=\langle e_{1},e_{2},\dots,e_{n}\rangle.$
Each event $e_{i}$ belongs to a finite set of observable event kinds, including:
  • I/O events: $\mathsf{Out}(s)$ and $\mathsf{Err}(s)$ for normalized writes to stdout/stderr.
  • Termination events: $\mathsf{Exit}(k)$ for a normal exit code $k$, and $\mathsf{Timeout}$ if the run exceeds the time budget.
  • Exception events: $\mathsf{Throw}(\tau,m)$ for uncaught exceptions, recording (a normalized form of) the exception type $\tau$ and message $m$.
  • Control-flow events: $\mathsf{Enter}(b)$ and $\mathsf{Leave}(b)$ when entering/leaving an instrumented basic block $b$.
  • Data-flow summaries: $\mathsf{State}(b,h)$ at selected boundaries, where $h$ is a checksum of in-scope variable values at block $b$.
  • Call events: $\mathsf{Call}(f,\vec{a})$ recording a normalized callee identifier $f$ and argument summary $\vec{a}$ for selected calls.
Normalization.
To make traces comparable across runs, OBsmith applies a deterministic normalization function $\mathsf{Norm}(\cdot)$ to event payloads (e.g., canonicalizing line endings, eliding runtime-specific prefixes in stack messages, and serializing values into a stable representation before hashing). In particular, $\mathsf{State}(b,h)$ uses a stable serializer and a hash function over primitive values and bounded summaries of objects/arrays to avoid brittleness due to formatting differences while still detecting semantic drift. We denote the resulting normalized trace by:
$\widehat{\mathsf{Trace}}_{E}(P,x)\triangleq\mathsf{Norm}\!\left(\mathsf{Trace}_{E}(P,x)\right).$
Operational semantic equivalence.
We define two programs $P$ and $Q$ to be observationally equivalent under environment $E$ and input $x$, written $P\approx_{E,x}Q$, iff their normalized traces are identical:
$P\approx_{E,x}Q\quad\Longleftrightarrow\quad\widehat{\mathsf{Trace}}_{E}(P,x)=\widehat{\mathsf{Trace}}_{E}(Q,x).$
This relation is intentionally stronger than “same final output”: it requires agreement on termination mode (normal exit vs. exception vs. timeout), the dynamic control-flow path (block enter/leave sequence), and a coarse-grained but effective approximation of data-flow and call behavior (state checksums and call-argument summaries).
Correctness criterion for obfuscation.
Let $\mathcal{O}$ be an obfuscator and $c$ an obfuscation configuration. OBsmith tests correctness by executing the original program $P$ and its obfuscated version $\mathcal{O}_{c}(P)$ on the same input $x$ and checking:
$P\approx_{E,x}\mathcal{O}_{c}(P).$
Any violation indicates a semantic divergence in the tested environment $E$ and under the operational criterion above.
Handling nondeterminism.
JavaScript programs and runtimes can exhibit nondeterminism (e.g., due to time, randomized hashing, or scheduling). OBsmith therefore (i) avoids nondeterministic APIs in generated tests by construction when possible, and (ii) reruns executions multiple times: a discrepancy is reported as a bug only if it reproduces consistently across reruns of both $P$ and $\mathcal{O}_{c}(P)$.
Limitations.
Our criterion does not claim full semantic equivalence for all JavaScript behaviors, especially those involving external I/O, the network, or environment-dependent APIs. Instead, it provides a practical, scalable, and substantially stronger-than-output notion of equivalence that is well-suited for detecting semantic drift introduced by source-to-source obfuscation.
3.7. Reference-Oriented Equivalence Testing
Differential testing is a widely adopted technique in compiler validation (Yang et al., 2011; Zang et al., 2023), where the absence of a reliable ground truth necessitates comparing the outputs of multiple compiler variants on the same input program. In contrast, OBsmith operates under the assumption that the original JavaScript program, executed directly in a trusted engine (Node.js), serves as the reference oracle (ground truth). This allows OBsmith to adapt reference-oriented equivalence testing specifically for the obfuscation domain.
@@ 3.7. Reference-Oriented Equivalence Testing
By leveraging the original program as a reference oracle, OBsmith strengthens differential testing with a reliable ground truth, termed reference-oriented equivalence testing. This adaptation enables precise detection of semantic inconsistencies introduced by obfuscation, distinguishing it from traditional approaches that rely on consensus among potentially flawed variants.
@@ 3.8. Metamorphic Testing
Reference-oriented equivalence testing (§3.7) is our primary mechanism for checking semantic equivalence between an original program and its obfuscated counterpart. In addition, we explore whether metamorphic testing can stress obfuscators by applying equivalence-preserving rewrites to the inputs before obfuscation and then verifying that obfuscation does not change program behavior. The goal here is not limited to the “single-obfuscator available” setting; rather, it is to probe whether obfuscators are robust to semantics-preserving variation in their inputs.
@@
OBsmith defines a suite of metamorphic relations (MRs) over JavaScript programs. Our implementation use three most frequently used MRs in compiler testing:
@@
Relations to obfuscation.
Metamorphic relations (MRs) provide test amplification rather than defining correctness. For an MR $m$, OBsmith first checks that $m$ preserves behavior on the original program, i.e., $P\approx_{E,x}m(P)$. Only then does it check that obfuscation preserves this equivalence as well, by comparing $\mathcal{O}_{c}(P)$ and $\mathcal{O}_{c}(m(P))$ against their respective references. This separates the system under test (the obfuscator) from the controlled transformations used to generate additional, independently-checkable test inputs.
@@
By encoding program equivalences as metamorphic relations, OBsmith provides a lightweight yet powerful mechanism to validate obfuscators in the absence of oracles. This technique, in combination with reference-oriented equivalence testing, yields comprehensive correctness checking across both multi-obfuscator and single-obfuscator scenarios
@@ 4.1.1. Large language models
Regarding LLM selection, we used the most capable LLMs available at the start of the experiment. For each LLM, we generate 100 different sketches. We also test the coding agent and see its effectiveness in generating sketchs compared to directly use base model. For fair comparison since coding agent like Gemini-CLI or Claude Code can see the whole project, we generate sketches first and then put other LLMs results into our project directory. Specifically, we use the following models: Gemini 2.5 Pro, GPT-5, Claude 4 Opus, Qwen3-235B-A22B-2507, Grok 4. Besides, we also evaluate LLM agent’s ability in sketch generation. We use Gemini CLI with Gemini 2.5 Pro to generate 100 sketches as well. In order to fair comparison, we begin with Gemini CLI without other models’ results, in other words, we do this first.
@@ 4.1.2. Obfuscators under test
Based on Google’s investigation (Jiang et al., 2026), Obfuscator.IO (obi, 2024) and JS-Confuser (jsc, 2022) are among the most widely used JavaScript obfuscators in industrial production. We therefore evaluate on these two tools to study failures in widely deployed obfuscation pipelines. We focus on two obfuscators to enable systematic coverage across configurations and to support thorough triage and root-cause analysis, rather than spreading the evaluation budget thinly across many tools.
Both tools take as input a candidate program and a JSON configuration that specifies obfuscation options. Because an obfuscator’s behavior can vary substantially with its settings, we run each candidate program under multiple configurations to generate a diverse set of obfuscated variants and increase coverage of the transformation space.
JS-Confuser provides three preset configurations (Low, Medium, High), and we use them without modification. Obfuscator.IO provides four presets (Default, Low, Medium, High), which we also use, with a small number of necessary adjustments for compatibility with our pipeline. Our implementation of OBsmith checks program equivalence using console.log; however, Obfuscator.IO’s Low/Medium/High presets disable the console, so we set the corresponding option to false. In addition, Obfuscator.IO’s High preset enables debug protection, which can cause scripts to hang forever, so we disable that option as well. Finally, we disable compact mode (which places the entire program on a single line), because V8 may reject extremely long lines with an “invalid size” error. We make no other configuration changes.